You purchase 8 gallons of paint and 3 brushes for $152.50.The next day, you purchase 6 gallons of paint and 2 brushes for $113.00. How much does each gallon of paint and each brush cost?
Answer by Paul(988)

Let paint be x and brushes be y
6x+2y=113
Solve for y
2y=113-6x
y=56.5-3x (subsitution)

8x+3y=152.5
Use subsitution:
8x+3(56.50-3x)=152.5
8x-9x=152.5-169.5
x=17

y=56.5-3(17)
y=5.5

Hence, one paint gallon costs $17 and one brush costs $5.5
Paul.